OpenStudy (anonymous):

13. You went to lunch with your friend, and the bill was $27.00. Your lunch is $8.00 more than your friend’s. How much was your friend’s lunch?

Directrix (directrix):

Let x be the cost of the friend's lunch. Then, your lunch is x + 8 So, x + x + 8 = 27 2x = 19 x = 19/2 = $9.50 for the friend. That doesn't look right, does it @123caca

Directrix (directrix):

Your lunch would be$ 17.50 17.50 + 9.50 = 27 So, I think this is correct for the friend. $9.50 for the friend. Do you, @123caca

OpenStudy (anonymous):

yes i do!! Thank you soooo much!!! @Directrix

Directrix (directrix):

You are welcome.